OpenStudy (anonymous):

The area of a rectangular park is 3/5 square mile .The length of the park is 7/8 miles .What is the width of the park. I lve you all so help me ill pray tonight for you fellow braniacs

OpenStudy (anonymous):

Area of a rectangle formula: A = L * W, where A is the area, L is the length, and W is the width. 3/5 = 7/8 * W Now, solve for W using algebra. (3/5) / (7/8) = (7/8 * W) / (7/8) (3/5) / (7/8) = W

OpenStudy (anonymous):

What's the slash sign mean

OpenStudy (anonymous):

The slashes are the same as the division sign. You can simplify from there. Remember that dividing by a fraction is the same as multiplying by the reciprocal of it (flip the numerator and denominator).

OpenStudy (anonymous):

I got an answer but can u simplify

OpenStudy (anonymous):

What answer did you get mrs

OpenStudy (anonymous):

Don't you multiplym 3/5 by 8/7 when you do the reciprocal

OpenStudy (anonymous):

\[\frac{3}{5} \div \frac{7}{8} = \frac{3}{5} \times \frac{8}{7} = \frac{24}{35}\] This fraction can't be simplified, as 24 and 35 don't have common factors. (By the way, I'm a guy :p Not offended or anything, it doesn't really matter)

OpenStudy (anonymous):

Ohh im really sorry